OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3. y − 6 = −3x − 3 y − 6 = −3(x + 1) y = −3x + 3 3x + y = 3

OpenStudy (ghazi):

use (y-y1)=m (x-x1) you'll get (y-6)= -3 (x+1)

OpenStudy (ghazi):

@Reemxoxo is this clear?